OpenStudy (desmarie):

F= 9/5(K−273) + 32 In order to convert a temperature in kelvin into a temperature in degrees in fahrenheit, the equation above can be used. Which of the following equations correctly expresses the temperature in kelvin in terms of the temperature in degrees Fahrenheit?

OpenStudy (desmarie):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`Which of the following equations correctly expresses the temperature in kelvin in terms of the temperature in degrees Fahrenheit?` translation: solve the given equation for K

OpenStudy (desmarie):

K=(5/9F - 32) + 273 ? ​

jimthompson5910 (jim_thompson5910):

Imagine we started with just K Let's apply a series of steps to get the right hand side of the equation step 1) subtract off 273 so we go from just K to K-273 step 2) multiply the result of step 1 by 9/5 we go from (K-273) to (9/5)*(K-273) step 3) Add 32 to the result of the prev step we go from (9/5)*(K-273) to (9/5)*(K-273)+32

OpenStudy (desmarie):

or would the 5/9 be on the outside of the parentheses??

jimthompson5910 (jim_thompson5910):

to recap so far, if we start with just K and go through step 1 through step 3 (shown above) then we end up at (9/5)*(K-273)+32

jimthompson5910 (jim_thompson5910):

to isolate K, we just undo everything going in reverse so we would start at step 3 and undo the addition of 32 which means we subtract 32 from both sides

jimthompson5910 (jim_thompson5910):

F= 9/5(K−273) + 32 F - 32= 9/5(K−273) + 32-32 ... undo step 3 (see above) F - 32 = 9/5(K−273) (5/9)*(F - 32) = 5/9*9/5(K−273) ... undo step 2 (see above) (5/9)*(F - 32) = K - 273 (5/9)*(F - 32)+273 = K - 273+273 ... undo step 1 (see above) (5/9)*(F - 32)+273 = K K = (5/9)*(F - 32)+273

OpenStudy (desmarie):

so the 5/9 has to be on the outside..

jimthompson5910 (jim_thompson5910):

yes applied to ALL of (F-32) not just the F

OpenStudy (desmarie):

THANK YOU x 100000000000000000

jimthompson5910 (jim_thompson5910):

you're welcome